Publications

G. A. Alexandrov, D. Ames, G. Bellocchi, M. Bruen, N. Crout, M. Erechtchoukova, A. Hildebrandt, F. Hoffman, C. Jackisch, P. Khaiter, G. Mannina, T. Matsunaga, S. T. Purucker, M. Rivington, and L. Samaniego. Technical assessment and evaluation of environmental models and software: Letter to the editor. Environ. Modell. Softw., 26(3):328–336, March 2011. doi:10.1016/j.envsoft.2010.08.004. Thematic issue on the assessment and evaluation of environmental models and software.

M. R. Allen, D. J. Erickson, R. J. Andres, F. M. Hoffman, and M. L. Branstetter. Monthly anthropogenic CO_{2} fluxes: Impacts on the atmospheric CO_{2} seasonal cycle and implications for models of the terrestrial biosphere. Abstract B41G-0391 presented at 2010 Fall Meeting, American Geophysical Union (AGU), San Francisco, California, USA, December 2010.

Barry Baker, William W. Hargrove, Forrest M. Hoffman, and Mike Heiner. Use of multivariate cluster and climate classification techniques to characterize future climate scenarios in Peoples Republic of China. In Proceedings of the 21st Annual Symposium of the International Association for Landscape Ecology, United States Regional Association (US-IALE), San Diego, California, USA, March 2006.

Barry Baker, Henry Diaz, William Hargrove, and Forrest Hoffman. Use of the Köppen-Trewartha climate classification to evaluate climatic refugia in statistically derived ecoregions for the People’s Republic of China. Clim. Change, 98(1):113–131, January 2010. ISSN 0165-0009. doi:10.1007/s10584-009-9622-2.

K. D. Barnes, J. M. Donato, D. M. Flanagan, N. W. Grady, J. A. Green, F. M. Hoffman, J. A. Kohl, M. R. Leuze, P. M. Papadopoulos, and R. F. Sincovec. The Financial automated Management On-line User System (FaMOUS): A prototype interactive hypertext-based financial planning and reporting system. Technical Memorandum ORNL/TM-13139, Oak Ridge National Laboratory, Oak Ridge, Tennessee, USA, November 1995.

William L. Bauerle, Ram Oren, Danielle A. Way, Song S. Qian, Paul C. Stoy, Peter E. Thornton, Joseph D. Bowden, Forrest M. Hoffman, and Robert F. Reynolds. Photoperiodic regulation of the seasonal pattern of photosynthetic capacity and the implications for carbon cycling. Proc. Nat. Acad. Sci., 109(22):8612–8617, May 2012. doi:10.1073/pnas.1119131109.

Bjørn J. Brooks, Forrest M. Hoffman, Richard T. Mills, David J. Erickson, and T. J. Blasing. The effect of anthropogenic emissions corrections on the seasonal cycle of atmospheric CO_{2}. Eos Trans. AGU, 90(52), December 2009. Fall Meet. Suppl., Abstract A51A-0108.

George R. Carr, Matthew J. Cordery, John B. Drake, Michael W. Ham, Forrest M. Hoffman, and Patrick H. Worley. Porting and performance of the Community Climate System Model (CCSM3) on the Cray X1. In Proceedings of the 2005 Cray Users Group (CUG) Conference, May 2005.

Robert E. Dickinson, Keith W. Oleson, Gordon Bonan, Forrest Hoffman, Peter Thornton, Mariana Vertenstein, Zong-Liang Yang, and Xubin Zeng. The Community Land Model and its climate statistics as a component of the Community Climate System Model. J. Clim., 19(11):2302–2324, June 2006. doi: 10.1175/JCLI3742.1.

Charles Ehlschlaeger, James Westervelt, Harold Balbach, H. Resit Akcakaya, Tom Hoctor, Crystal Goodison, William W. Hargrove, Forrest M. Hoffman, Winifred Rose, and Robert C. Lozar. Habitat fragmentation handbook for installation planners. Technical Report ERDC/CERL TR-06-36, U.S. Army Corps of Engineers, Engineer Research and Development Center, December 2006.

David Erickson, Robert Oglesby, Scott Elliott, and Forrest Hoffman. Peta-scale climate modeling: Biogeochemical and financial feedbacks. In Proceedings of the iEMSs Third Biennial Meeting: International Congress on Environmental Modelling and Software Society (iEMSs 2006), Summit on Environmental Modelling and Software, Burlington, Vermont, USA, July 2006a.

David J. Erickson, T. J. Blasing, Forrest M. Hoffman, Richard T. Mills, Z. Zhu, and Stephan R. Kawa. Monthly global emissions of anthropogenic CO_{2}: Atmospheric CO_{2} transport calculations based on NASA data assimilation. Eos Trans. AGU, 87(52), December 2006b. Fall Meet. Suppl., Abstract A41C-0044.

David J. Erickson III, Richard T. Mills, Jay Gregg, T. J. Blasing, Forrest M. Hoffman, Robert J. Andres, Matthew Devries, Z. Zhu, and S. R. Kawa. An estimate of monthly global emissions of anthropogenic CO_{2}: Impact on the seasonal cycle of atmospheric CO_{2}. J. Geophys. Res., 113(G1):G01023, March 2008. doi:10.1029/2007JG000435.

David J. Erickson, Auroop Ganguly, Karsten Steinhaeuser, Marcia Branstetter, Robert J. Oglesby, Forrest M. Hoffman, and Lawrence Buja. Extreme climate event trends: The data mining and evaluation of the A1FI scenario for 2000–2100. Eos Trans. AGU, 89(53), December 2008. Fall Meet. Suppl., Abstract H12B-03, Invited.

David J. Erickson, Steven Pawson, Jamison Daniel, Melissa Allen, L. E. Ott, Auroop Ganguly, E. Nielsen, and Forrest M. Hoffman. Atmospheric CO_{2} simulation inside GEOS-5: Data mining, evaluation and treaty verification. Eos Trans. AGU, 90(52), December 2009. Fall Meet. Suppl., Abstract A51A-0078.

Jin-Ping Gwo, Forrest M. Hoffman, and William W. Hargrove. Mechanistic-based genetic algorithm search on a Beowulf cluster of Linux PCs. In Proceedings of the High Performance Computing 2000 (HPC2000) Conference, April 2000.

Jin-Ping Gwo, Eduardo F. D’Azevedo, Hartmut Frenzel, Melaine Mayes, Gour-Tsyh Yeh, Philip M. Jardine, Karen M. Salvage, and Forrest M. Hoffman. HBGC123D: A high performance computer model of coupled hydrogeological and biogeochemical processes. Comput. Geosci., 27(10):1231–1242, December 2001. doi: 10.1016/S0098-3004(01)00027-9.

William W. Hargrove, Forrest M. Hoffman, and Daniel A. Levine. Interpolation of bottom bathymetry and potential erosion in a large Tennessee reservoir system using GRASS. In Proceedings of the Ninth Annual Symposium on Geographic Information Systems, pages 552–557, March 1995.

William W. Hargrove and Forrest M. Hoffman. Using multivariate clustering to characterize ecoregion borders. Comput. Sci. Eng., 1(4):18–25, July 1999. doi:10.1109/5992.774837.

William W. Hargrove and Forrest M. Hoffman. An analytical assessment tool for predicting changes in a species distribution map following changes in environmental conditions. In B. O. Parks, K. M. Clarke, and M. P. Crane, editors, Proceedings of the Fourth International Conference on Integrating GIS and Environmental Modeling (GIS/EM4): Problems, Prospects and Research Needs, Boulder, Colorado, September 2000. University of Colorado, Cooperative Institute for Research in Environmental Sciences (CIRES).

ISBN 0-9743307-0-1. URL http://www.colorado.edu/research/cires/banff/pubpapers/104/.

William W. Hargrove, Forrest M. Hoffman, and Thomas Sterling. The do-

it-yourself supercomputer.  Sci. Am., 265(2):72–79, August 2001.  URL

http://www.sciam.com/article.cfm?articleID=000E238B-33EC-1C6F-84A9809EC588EF21.

William W. Hargrove, Forrest M. Hoffman, and Paul M. Schwartz. A fractal landscape realizer for generating synthetic maps. Conserv. Ecol., 6(1):2, February 2002. URL http://www.consecol.org/vol6/iss1/art2/. Part of Special Feature on Ralf Yorque Memorial Competition 2001.

William W. Hargrove, Forrest M. Hoffman, and Beverly E. Law. New analysis reveals representativeness of the AmeriFlux Network. Eos Trans. AGU, 84(48):529, 535, December 2003. doi:10.1029/2003EO480001.

William W. Hargrove and Forrest M. Hoffman. Developing a practical map-analysis tool for corridor detection. In Proceedings of the 19th Annual Symposium of the International Association for Landscape Ecology, United States Regional Association (US-IALE), pages 93–94, Las Vegas, Nevada, March 2004a.

William W. Hargrove and Forrest M. Hoffman. A flux atlas for representativeness and statistical extrapolation of the AmeriFlux network. Technical Memorandum ORNL/TM-2004/112, Oak Ridge National Laboratory, Oak Ridge, Tennessee, USA, April 2004b. URL http://www.geobabble.org/flux-ecoregions/.

William W. Hargrove and Forrest M. Hoffman. Potential of multivariate quantitative methods for delineation and visualization of ecoregions. Environ. Manage., 34(Supplement 1):S39–S60, April 2004c. doi:10.1007/s00267-003-1084-0.

William W. Hargrove, Forrest M. Hoffman, and Rebecca A. Efroymson. A practical map-analysis tool for detecting potential dispersal corridors. Landscape Ecol., 20(4):361–373, May 2005a. doi: 10.1007/s10980-004-3162-y.

William W. Hargrove and Forrest M. Hoffman. Quantifying representativeness importance values for AmeriFlux sites. Eos Trans. AGU, 86(52), December 2005. Fall Meet. Suppl., Abstract B51C-0219.

William W. Hargrove, Forrest M. Hoffman, Bruce P. Hayden, Dean L. Urban, J. A. MacMahon, and J. F. Franklin. Development of a domain map for nodes of the National Ecological Observatory Network (NEON). Eos Trans. AGU, 86(52), December 2005b. Fall Meet. Suppl., Abstract H33H-02.

William W. Hargrove and Forrest M. Hoffman. Quantifying representativeness importance values for ameriflux tower locations. In Anni Reissel and Asbjørn Aarflot, editors, Proceedings of the 1st Integrated Land Ecosystem-Atmsophere Processes Study (iLEAPS) Science Conference, Boulder, Colorado, number 79 in Report Series in Aerosol Science, page 252, Helsinki, Finland, January 2006a. ISBN 952-5027-66-X. ISSN 0784-3496.

William W. Hargrove, Bruce Hayden, Dean Urban, James MacMahon, Jerry Franklin, and Forrest M. Hoffman. Development of a domain map for nodes of the National Ecological Observatory Network (NEON). In Proceedings of the 21st Annual Symposium of the International Association for Landscape Ecology, United States Regional Association (US-IALE), San Diego, California, USA, March 2006a.

William W. Hargrove, Forrest M. Hoffman, and Paul F. Hessburg. Mapcurves: A quantitative method for comparing categorical maps. J. Geograph. Syst., 8(2):187–208, July 2006b. doi:10.1007/s10109-006-0025-x.

William W. Hargrove and Forrest M. Hoffman. Multivariate geographic clustering as a basis for ecoregionalization in the environmental sciences. Eos Trans. AGU, 87(52), December 2006b. Fall Meet. Suppl., Abstract IN41C-02, Invited.

William W. Hargrove, Joe Spruce, Gerry Gasser, Forrest M. Hoffman, and Danny Lee. A new national MODIS-derived phenology data set every 16 days, 2002 through 2006. Eos Trans. AGU, 89(53), December 2008. Fall Meet. Suppl., Abstract B51B-0373.

William W. Hargrove, Joseph P. Spruce, Gerald E. Gasser, and Forrest M. Hoffman. Toward a national early warning system for forest disturbances using remotely sensed phenology. Photogramm. Eng. Rem. Sens., 75(10):1150–1156, October 2009a.

William W. Hargrove, Joseph P. Spruce, Gerald E. Gasser, and Forrest M. Hoffman. Toward a national early warning system for forest disturbances using remotely sensed land surface phenology. Eos Trans. AGU, 90(52), December 2009b. Fall Meet. Suppl., Abstract B42B-06.

Paul Hessburg, William W. Hargrove, and Forrest M. Hoffman. A quantitative method for comparing categorical maps. In Proceedings of the 21st Annual Symposium of the International Association for Landscape Ecology, United States Regional Association (US-IALE), San Diego, California, USA, March 2006.

Forrest M. Hoffman. A Unix print filter for controlling an HP Laserjet printer. Technical Memorandum ORNL/TM-12190, Oak Ridge National Laboratory, Oak Ridge, Tennessee, USA, September 1992.

Forrest M. Hoffman and Vijay S. Tripathi. A geochemical expert system prototype using object-oriented knowledge representation and a production rule system. Comput. Geosci., 19(1):53–60, January 1993. doi:10.1016/0098-3004(93)90042-4.

Forrest M. Hoffman. How do I connect to the Internet? let me count the ways. In Proceedings of WATTec ’94, Knoxville, Tennessee, February 1994a.

Forrest M. Hoffman. Converting hard copy document for electronic dissemination. In Proceedings of the 18th Annual Practical Conference on Communication (PCOC), Oak Ridge, Tennessee, November 1994b.

Forrest M. Hoffman. Deployment of Internet technologies at Oak Ridge National Laboratory. In Proceedings of WATTec ’95, Knoxville, Tennessee, February 1995.

Forrest M. Hoffman and William W. Hargrove. Making soup from stones. Troubleshooting Professional, 2 (5), May 1998.

Forrest M. Hoffman and William W. Hargrove. Multivariate geographic clustering using a Beowulf-style parallel computer. In Hamid R. Arabnia, editor, Proceedings of the International Conference on Parallel and Distributed Processing Techniques and Applications (PDPTA ’99), volume III, pages 1292–1298. CSREA Press, June 1999a. ISBN 1-892512-11-4.

Forrest M. Hoffman and William W. Hargrove. Parallel computing with Linux. Crossroads, 6(1):23–27, September 1999b. doi:10.1145/331636.331643.

Forrest M. Hoffman and William W. Hargrove. High performance computing: An introduction to parallel programming with Beowulf. Open Source Developers Journal, 1(1):24–31, 2000.

Forrest M. Hoffman, Robert J. Oglesby, William W. Hargrove, and David J. Erickson. Using clustering to establish climate regimes from PCM output. Eos Trans. AGU, 83(47), December 2002. Fall Meet. Suppl., Abstract A61C-0090.

Forrest M. Hoffman, William W. Hargrove, David J. Erickson, and Robert J. Oglesby. Using clustered climate regimes for understanding water cycle variability. In Proceedings of the American Meteorological Society 83rd Annual Meeting, volume 69, Long Beach, California, February 2003a. Abstract 9.4.

Forrest M. Hoffman, William W. Hargrove, and Anthony D. Del Genio. Multivariate spatio-temporal clustering of time-series data: An approach for diagnosing cloud properties and understanding ARM site representativeness. In Proceedings of the U.S. Department of Energy Atmospheric Radiation Measurement (ARM) Program Science Team Meeting, Broomfield, Colorado, April 2003b. Extended Abstract.

Forrest M. Hoffman, William W. Hargrove, David J. Erickson, and Robert J. Oglesby. A novel method for analyzing and interpreting GCM results using clustered climate regimes. Eos Trans. AGU, 84(46), December 2003c. Fall Meet. Suppl., Abstract GC12A-0155.

Forrest M. Hoffman, Mariana Vertenstein, Hideyuki Kitabata, James B. White, Patrick Worley, John Drake, and Matthew Cordery. Adventures in vectorizing the Community Land Model. In Proceedings of the 2004 Cray Users Group (CUG) Conference, May 2004a.

Forrest Hoffman, Mariana Vertenstein, Peter Thornton, Keith Oleson, and Samuel Levis. Community Land Model version 3.0 (CLM3.0) developer’s guide. Technical Memorandum ORNL/TM-2004/119, Oak Ridge National Laboratory, Oak Ridge, Tennessee, USA, June 2004b. URL

http://www.cgd.ucar.edu/tss/clm/distribution/clm3.0/DevelopersGuide/doc/CodeReference/DevGuideAndReference.pdf.

Forrest M. Hoffman. Analysis of reflected spectral signatures and detection of geophysical disturbance using hyperspectral imagery. Master’s thesis, University of Tennessee, Department of Physics and Astronomy, Knoxville, Tennessee, USA, November 2004.

Forrest M. Hoffman and William W. Hargrove. Quantitative flux ecoregions for AmeriFlux using MODIS. Eos Trans. AGU, 85(47), December 2004. Fall Meet. Suppl., Abstract B23B-07.

Forrest M. Hoffman, William W. Hargrove, David J. Erickson, and Robert J. Oglesby. Using clustered climate regimes to analyze and compare predictions from fully coupled general circulation models. Earth Interact., 9(10):1–27, August 2005a. doi:10.1175/EI110.1.

Forrest M. Hoffman, Mariana Vertenstein, Hideyuki Kitabata, and James B. White III. Vectorizing the Community Land Model (CLM). Int. J. High Perf. Comput. Appl., 19(3):247–260, August 2005b. doi: 10.1177/1094342005056113.

Forrest M. Hoffman, Inez Fung, and Jasmin John. Preliminary results from the C^{4}MIP Phase 1 simulations using the CCSM3-CLM3-CASA′ coupled model. Eos Trans. AGU, 86(52), December 2005c. Fall Meet. Suppl., Abstract B33G-07.

Forrest M. Hoffman, Peter Thornton, Inez Fung, and W. Mac Post. Land-atmosphere interactions exhibited by coupled carbon-cycle climate models. In Anni Reissel and Asbjørn Aarflot, editors, Proceedings of the 1st Integrated Land Ecosystem-Atmsophere Processes Study (iLEAPS) Science Conference, Boulder, Colorado, number 79 in Report Series in Aerosol Science, page 341, Helsinki, Finland, January 2006a. ISBN 952-5027-66-X. ISSN 0784-3496.

Forrest M. Hoffman and William W. Hargrove. Applying quantitative ecoregionalization to network analysis: Quantifying representativeness and determining importance values for AmeriFlux sites. In Proceedings of the 21st Annual Symposium of the International Association for Landscape Ecology, United States Regional Association (US-IALE), San Diego, California, USA, March 2006.

Forrest M. Hoffman, Inez Fung, W. Mac Post, and David Erickson. Recent results from coupled climate/carbon-cycle models in CCSM3. In Proceedings of the iEMSs Third Biennial Meeting: International Congress on Environmental Modelling and Software Society (iEMSs 2006), Summit on Environmental Modelling and Software, Burlington, Vermont, USA, July 2006b.

Forrest M. Hoffman, Inez Fung, Jim Randerson, Peter Thornton, Jon Foley, Curtis Covey, Jasmin John, Samuel Levis, W. Mac Post, Mariana Vertenstein, Reto Stöckli, Steve Running, Faith Ann Heinsch, David Erickson, and John Drake. Terrestrial biogeochemistry in the Community Climate System Model (CCSM). J. Phys.: Conf. Ser., 46(1):363–369, September 2006c. doi:10.1088/1742-6596/46/1/051.

Forrest M. Hoffman, Inez Y. Fung, James T. Randerson, Peter E. Thornton, Reto Stöckli, Faith Ann Heinsch, Steve Running, Kathy Hibbard, Jasmin John, Curt Covey, Jon Foley, W. Mac Post, William W. Hargrove, David J. Erickson, and Natalie Mahowald. Preliminary results from the CCSM Carbon-Land Model Intercomparison Project (C-LAMP). Eos Trans. AGU, 87(52), December 2006d. Fall Meet. Suppl., Abstract B51C-0316.

Forrest M. Hoffman, Curtis C. Covey, Inez Y. Fung, James T. Randerson, Peter E. Thornton, Yen-Huei Lee, Nan A. Rosenbloom, Reto C. Stöckli, Steven W. Running, David E. Bernholdt, and Dean N. Williams. Results from the Carbon-Land Model Intercomparison Project (C-LAMP) and availability of the data on the Earth System Grid (ESG). J. Phys.: Conf. Ser., 78(1):012026, December 2007a. doi:10.1088/1742-6596/78/1/012026.

Forrest M. Hoffman, James T. Randerson, Inez Y. Fung, Peter E. Thornton, Jeff Lee, and Curt Covey. Results from the Carbon-Land Model Intercomparison Project (C-LAMP). Eos Trans. AGU, 88(52), December 2007b. Fall Meet. Suppl., Abstract B31C-0324.

Forrest M. Hoffman, William W. Hargrove, Richard T. Mills, Salil Mahajan, David J. Erickson, and Robert J. Oglesby. Multivariate Spatio-Temporal Clustering (MSTC) as a data mining tool for environmental applications. In Miquel Sànchez-Marrè, Javier Béjar, Joaquim Comas, Andrea E. Rizzoli, and Giorgio Guariso, editors, Proceedings of the iEMSs Fourth Biennial Meeting: International Congress on Environmental Modelling and Software Society (iEMSs 2008), pages 1774–1781, July 2008a. ISBN 978-84-7653-074-0.

Forrest M. Hoffman, James T. Randerson, Inez Y. Fung, Peter E. Thornton, Yen-Huei “Jeff” Lee, Curtis C. Covey, Gordon B. Bonan, and Steven W. Running. The Carbon-Land Model Intercomparison Project (C-LAMP): A protocol and evaluation metrics for global terrestrial biogeochemistry models. In Miquel Sànchez-Marrè, Javier Béjar, Joaquim Comas, Andrea E. Rizzoli, and Giorgio Guariso, editors, Proceedings of the iEMSs Fourth Biennial Meeting: International Congress on Environmental Modelling and Software Society (iEMSs 2008), pages 1039–1046, July 2008b. ISBN 978-84-7653-074-0.

Forrest M. Hoffman, James T. Randerson, Inez Y. Fung, Peter Thornton, Curtis Covey, Gordon Bonan, Steven Running, and Richard Norby. Comparison of global model results from the Carbon-Land Model Intercomparison Project (C-LAMP) with Free-Air Carbon Dioxide Enrichment (FACE) manipulation experiments. Eos Trans. AGU, 89(53), December 2008c. Fall Meet. Suppl., Abstract B51E-0447.

Forrest M. Hoffman and Martial Mancip. Working group report on terrestrial biosphere model evaluation. Integrated Land Ecosystem-Atmosphere Processes Study (iLEAPS) Newsletter, 7:64, June 2009. ISSN 1796-0363.

Forrest M. Hoffman, James T. Randerson, Inez Y. Fung, Peter E. Thornton, Natalie M. Mahowald, Gordon B. Bonan, and Steven W. Running. The Carbon-Land Model Intercomparison Project (C-LAMP): A prototype for coupled biosphere-atmosphere model benchmarking for the IPCC Fifth Assessment Report (AR5). In Anni Reissell, Marjut Nyman, Miia Vesala, and Tyyne Viisanen, editors, Water in a Changing Climate — Progress in Land-Atmosphere Interactions and Energy/Water Cycle Research, volume 1 of Proceedings of the 6th International Scientific Conference on the Global Energy and Water Cycle (GEWEX) and 2nd International Land Ecosystem-Atmosphere Processes Study (iLEAPS) Science Conference, pages 126–127, Melbourne, Australia, August 2009a. ISBN 978-952-5855-01-2.

Forrest M. Hoffman, James T. Randerson, Peter E. Thornton, Natalie M. Mahowald, Gordon B. Bonan, Steven W. Running, and Inez Y. Fung. The Carbon-Land Model Intercomparison Project (C-LAMP): A model-data comparison system for evaluation of coupled biosphere-atmosphere models. In Proceedings of the 8th International Carbon Dioxide Conference, Jena, Germany, September 2009b.

Forrest M. Hoffman, James T. Randerson, Peter E. Thornton, Gordon B. Bonan, Bjørn J. Brooks, and Inez Y. Fung. An international land-biosphere model benchmarking activity for the IPCC Fifth Assessment Report (AR5). Eos Trans. AGU, 90(52), December 2009c. Fall Meet. Suppl., Abstract B23G-08.

Forrest M. Hoffman, Richard T. Mills, Jitendra Kumar, Srinivasa S. Vulli, and William W. Hargrove. Geospatiotemporal data mining in an early warning system for forest threats in the United States. In Proceedings of the 2010 IEEE International Geoscience and Remote Sensing Symposium (IGARSS 2010), pages 170–173, July 2010. ISBN 978-1-4244-9566-5. doi:10.1109/IGARSS.2010.5653935. Invited.

F. M. Hoffman and J. T. Randerson. The impact of the temperature sensitivity of ecosystem respiration on the climate-carbon cycle feedback strength. Abstract B44A-02 presented at 2010 Fall Meeting, American Geophysical Union (AGU), San Francisco, California, USA, December 2010.

Forrest M. Hoffman, J. Walter Larson, Richard Tran Mills, Bjørn-Gustaf J. Brooks, Auroop R. Ganguly, William W. Hargrove, Jian Huang, Jitendra Kumar, and Ranga R. Vatsavai. Data Mining in Earth System Science (DMESS 2011). In Mitsuhisa Sato, Satoshi Matsuoka, Peter M. Sloot, G. Dick van Albada, and Jack Dongarra, editors, Proceedings of the International Conference on Computational Science (ICCS 2011), volume 4 of Procedia Comput. Sci., pages 1450–1455, Amsterdam, June 2011. Elsevier. doi: 10.1016/j.procs.2011.04.157.

D. N. Huntzinger, W. M. Post, Y. Wei, A. M. Michalak, T. O. West, A. R. Jacobson, I. T. Baker, J. M. Chen, K. J. Davis, D. J. Hayes, F. M. Hoffman, A. K. Jain, S. Liu, A. D. McGuire, R. P. Neilson, Chris Potter, B. Poulter, David Price, B. M. Raczka, H. Q. Tian, P. Thornton, E. Tomelleri, N. Viovy, J. Xiao, W. Yuan, N. Zeng, M. Zhao, and R. Cook. North American Carbon Program (NACP) regional interim synthesis: Terrestrial biospheric model intercomparison. Ecol. Model., 232(0):144–157, May 2012. doi:10.1016/j.ecolmodel.2012.02.004.

C. Ryan Johnson, Markus Glatter, Wesley Kendall, Jian Huang, and Forrest M. Hoffman. Querying for feature extraction and visualization in climate modeling. In Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M.A. Sloot, editors, Proceedings of the 9th International Conference on Computational Science (ICCS 2009), volume 5545 of Lecture Notes in Computer Science (LNCS), pages 416–425, Heidelberg, May 2009. Springer-Verlag. ISBN 978-3-642-01972-2. doi:10.1007/978-3-642-01973-9 46.

Michael Keller, David Schimel, William Hargrove, and Forrest Hoffman. A continental strategy for the National Ecological Observatory Network. Front. Ecol. Environ., 6(5):282–284, June 2008. doi: 10.1890/1540-9295(2008)6[282:ACSFTN]2.0.CO;2. Special Issue on Continental-Scale Ecology.

Wesley Kendall, Markus Glatter, Jian Huang, Forrest Hoffman, and David E. Bernholdt. Web enabled collaborative climate visualization in the earth system grid. In Proceedings of the International Symposium on Collaborative Technologies and Systems 2008 (CTS 2008), pages 212–220, May 2008. ISBN 978-1-4244-2248-7. doi:10.1109/CTS.2008.4543934.

Silvia Kloster, Natalie M. Mahowald, James T. Randerson, Peter E. Thornton, Forrest M. Hoffman, Samuel Levis, Peter J. Lawrence, Johan J. Feddema, Keith W. Oleson, and David M. Lawrence. Fire dynamics during the 20th century simulated by the Community Land Model. Biogeosci., 7(6):1877–1902, June 2010. doi:10.5194/bg-7-1877-2010.

Jitendra Kumar, Richard Tran Mills, Forrest M. Hoffman, and William W. Hargrove. Parallel k-means clustering for quantitative ecoregion delineation using large data sets. In Mitsuhisa Sato, Satoshi Matsuoka, Peter M. Sloot, G. Dick van Albada, and Jack Dongarra, editors, Proceedings of the International Conference on Computational Science (ICCS 2011), volume 4 of Procedia Comput. Sci., pages 1602–1611, Amsterdam, June 2011. Elsevier. doi:10.1016/j.procs.2011.04.173.

S. Y. Lee, Mark Elless, and Forrest M. Hoffman. Solubility measurement of Uranium in Uraniumcontaminated soils. Technical Memorandum ORNL/TM-12401, Oak Ridge National Laboratory, Oak Ridge, Tennessee, USA, August 1993.

Daniel A. Levine, William W. Hargrove, and Forrest M. Hoffman. Characterization of sediments in the Clinch River, Tennessee, using remote sensing and multi-dimensional GIS techniques. In Proceedings of the Ninth Annual Symposium on Geographic Information Systems, pages 548–551, March 1995a.

Daniel A. Levine, William W. Hargrove, and Forrest M. Hoffman. Characterization of sediments in the Clinch River, Tennessee, using remote sensing and multi-dimensional GIS techniques. In Michael Heit, H. Dennison Parker, and Art Shortreid, editors, GIS Applications in Natural Resources 2. GIS World, Inc., Fort Collins, Colorado, August 1995b. ISBN 1-882610-17-2.

Robert C. Lozar, William Hargrove, and Forrest Hoffman. Use of the Corridor Tool in support of threatened and endangered species habitat fragmentation. Technical Report ERDC/CERL TR-05-23, U.S. Army Corps of Engineers, Engineer Research and Development Center, September 2005.

Salil Mahajan, Forrest M. Hoffman, William W. Hargrove, Sigurd W. Christensen, and Richard T. Mills. A cluster analysis approach to comparing Atmospheric Radiation Measurement (ARM) observations with global climate model (GCM) results. Eos Trans. AGU, 88(52), December 2007. Fall Meet. Suppl., Abstract A41A-0010.

Gnanamanika Mahinthakumar, Forrest M. Hoffman, William W. Hargrove, and Nicolas T. Karonis. Multivariate geographic clustering in a metacomputing environment using Globus. In Supercomputing ’99: Proceedings of the 1999 ACM/IEEE conference on Supercomputing (CDROM), Supercomputing ’99, New York, NY, USA, November 1999. ACM Press. ISBN 1-58113-091-0. doi:10.1145/331532.331537.

N. M. Mahowald, S. Kloster, S. Engelstaedter, J. K. Moore, S. Mukhopadhyay, J. R. McConnell, S. Albani, S. C. Doney, A. Bhattacharya, M. A. J. Curran, M. G. Flanner, F. M. Hoffman, D. M. Lawrence, K. Lindsay, P. A. Mayewski, J. Neff, D. Rothenberg, E. Thomas, P. E. Thornton, and C. S. Zender. Observed 20th century desert dust variability: Impact on climate and biogeochemistry. Atmos. Chem. Phys., 10 (22):10875–10893, November 2010a. doi:10.5194/acp-10-10875-2010.

N. M. Mahowald, S. Kloster, S. Engelstaedter, J. K. Moore, S. Mukhopadhyay, J. R. McConnell, S. Albani, S. C. Doney, A. Bhattacharya, M. A. J. Curran, M. G. Flanner, F. M. Hoffman, D. M. Lawrence, K. Lindsay, P. A. Mayewski, J. Neff, D. Rothenberg, E. Thomas, and P. E. Thornton. Observed 20th century desert dust variability: Impact on climate and biogeochemistry. Abstract A11K-02 presented at 2010 Fall Meeting, American Geophysical Union (AGU), San Francisco, California, USA, December 2010b.

Richard T. Mills, Forrest M. Hoffman, Patrick H. Worley, Kalyan S. Perumalla, Art Mirin, Glenn E. Hammond, and Barry F. Smith. Coping at the user-level with resource limitations in the Cray Message Passing Toolkit MPI at scale: How not to spend your summer vacation. In Proceedings of the 2009 Cray User Group (CUG) Conference, May 2009.

R. T. Mills, F. M. Hoffman, J. Kumar, S. S. Vulli, W. W. Hargrove, and J. Spruce. Geospatiotemporal data mining of remotely sensed phenology for unsupervised forest threat detection. Abstract B23G-0462 presented at 2010 Fall Meeting, American Geophysical Union (AGU), San Francisco, California, USA, December 2010.

Richard Tran Mills, Forrest M. Hoffman, Jitendra Kumar, and William W. Hargrove. Cluster analysis-based approaches for geospatiotemporal data mining of massive data sets for identification of forest threats. In Mitsuhisa Sato, Satoshi Matsuoka, Peter M. Sloot, G. Dick van Albada, and Jack Dongarra, editors, Proceedings of the International Conference on Computational Science (ICCS 2011), volume 4 of Procedia Comput. Sci., pages 1612–1621, Amsterdam, June 2011. Elsevier. doi:10.1016/j.procs.2011.04.174.

Richard J. Norby, J. M. Warren, C. M. Iversen, B. E. Medlyn, R. E. McMurtrie, and Forrest M. Hoffman. Nitrogen limitation is reducing the enhancement of NPP by elevated CO_{2} in a deciduous forest. Eos Trans. AGU, 89(53), December 2008. Fall Meet. Suppl., Abstract B32B-05, Invited.

Keith W. Oleson, David M. Lawrence, Gordon B. Bonan, Mark G. Flanner, Erik Kluzek, Peter J. Lawrence, Samuel Levis, Sean C. Swenson, Peter E. Thornton, Aiguo Dai, Mark Decker, Robert Dickinson, Johannes Feddema, Colette Heald, Forrest Hoffman, Jean-François Lamarque, Natalie Mahowald, Guo-Yue Niu, Taotao Qian, James Randerson, Steve Running, Koichi Sakaguchi, Andrew Slater, Reto Stöckli, Aihui Wang, Zong-Liang Yang, Xiaodong Zeng, and Xubin Zeng. Technical description of version 4.0 of the Community Land Model (CLM). Technical Note NCAR/TN-478+STR, National Center for Atmospheric Research, Boulder, Colorado, USA, March 2010. URL

http://www.cesm.ucar.edu/models/cesm1.0/clm/CLM4_Tech_Note.pdf.

Keith W. Oleson, Yongjiu Dai, Gordon Bonan, Mike Bosilovich, Robert Dickinson, Paul Dirmeyer, Forrest Hoffman, Paul Houser, Samuel Levis, Guo-Yue Niu, Peter Thornton, Mariana Vertenstein, Zong-Liang Yange, and Xubin Zeng. Technical description of the Community Land Model. Technical Note NCAR/TN-461+STR, National Center for Atmospheric Research, Boulder, Colorado, USA, May 2004.

URL http://www.cgd.ucar.edu/tss/clm/distribution/clm3.0/TechNote/CLM_Tech_Note.pdf.

Jasna V. Pittman, Elliot M. Weinstock, Robert J. Oglesby, David S. Sayres, Jessica B. Smith, James G. Anderson, Owen R. Cooper, Steven C. Wofsy, Irene Xueref, Cristoph Gerbig, Bruce C. Daube, Erik C. Richard, Brian A. Ridley, Andrew J. Weinheimer, Max Loewenstein, Hans-Jurg Jost, Jimena P. Lopez, Michael J. Mahoney, Thomas L. Thompson, William W. Hargrove, and Forrest M. Hoffman. Transport in the subtropical lowermost stratosphere during the Cirrus Regional Study of Tropical Anvils and Cirrus Layers-Florida Area Cirrus Experiment. J. Geophys. Res., 112(D8):D08304, April 2007. doi: 10.1029/2006JD007851.

Wilfred M. Post, Anthony W. King, Stan D. Wullschleger, and Forrest M. Hoffman. Historical variations in terrestrial biospheric carbon storage. DOE Research Summary, (34), June 1997. URL

http://cdiac.esd.ornl.gov/pns/doers/doer34/doer34.htm.

James T. Randerson, Forrest M. Hoffman, Peter E. Thornton, Natalie M. Mahowald, Keith Lindsay, Yen-Huei Lee, Cynthia D. Nevison, Scott C. Doney, Gordon Bonan, Reto Stöckli, Curtis Covey, Steven W. Running, and Inez Y. Fung. Systematic assessment of terrestrial biogeochemistry in coupled climate-carbon models. Global Change Biol., 15(10):2462–2484, October 2009. ISSN 1365-2486. doi:10.1111/j.1365-2486.2009.01912.x.

Earl Saxon, Barry Baker, William Hargrove, Forrest Hoffman, and Chris Zganjar. Mapping environments at risk under different global climate change scenarios. Ecol. Lett., 8(1):53–60, January 2005. doi: 10.1111/j.1461-0248.2004.00694.x.

David Schimel, William Hargrove, Forrest Hoffman, and James McMahon. NEON: A hierarchically designed national ecological network. Front. Ecol. Environ., 5(2):59, March 2007. doi: 10.1890/1540-9295(2007)5[59:NAHDNE]2.0.CO;2.

X. Shi, J. Mao, P. E. Thornton, and F. M. Hoffman. The impact of climate, CO_{2}, nitrogen deposition and land use change on contemporary global river flow. Abstract B41G-0401 presented at 2010 Fall Meeting, American Geophysical Union (AGU), San Francisco, California, USA, December 2010.

Robert Sisneros, Markus Glatter, Brandon Langley, Jian Huang, Forrest Hoffman, and David J. Erickson III. Time-varying multivariate visualization for understanding terrestrial biogeochemistry. J. Phys.: Conf. Ser., 125(1):012093, December 2008. doi:10.1088/1742-6596/125/1/012093.

Robert Sisneros, Jian Huang, George Ostrouchov, and Forrest Hoffman. Visualizing life zone boundary sensitivities across climate models and temporal spans. In Mitsuhisa Sato, Satoshi Matsuoka, Peter M. Sloot, G. Dick van Albada, and Jack Dongarra, editors, Proceedings of the International Conference on Computational Science (ICCS 2011), volume 4 of Procedia Comput. Sci., pages 1582–1591, Amsterdam, June 2011. Elsevier. doi:10.1016/j.procs.2011.04.171.

Peter E. Thornton, Forrest M. Hoffman, and George C. Hurtt. Influence of dynamic land use and land cover change on simulated global terrestrial carbon and nitrogen cycles, climate-carbon cycle feedbacks, and interactions with rising CO_{2} and anthropogenic nitrogen deposition. Eos Trans. AGU, 90(52), December 2009. Fall Meet. Suppl., Abstract B24B-06.

Michael A. Unseren and Forrest M. Hoffman. Errata report on Herbert Goldstein’s Classical Mechanics, second edition. Technical Memorandum ORNL/TM-12176, Oak Ridge National Laboratory, Oak Ridge, Tennessee, USA, January 1993.

Mariana Vertenstein, Keith Oleson, Sam Levis, and Forrest Hoffman. Community Land Model version 3.0 (CLM3.0) user’s guide. Technical report, National Center for Atmospheric Research, Boulder, Colorado, USA, June 2004. URL

http://www.cgd.ucar.edu/tss/clm/distribution/clm3.0/UsersGuide/UsersGuide.pdf.

Michael A. White, Forrest M. Hoffman, and William W. Hargrove. A strategy for global phenological observatories. Eos Trans. AGU, 85(47), December 2004. Fall Meet. Suppl., Abstract B44A-02.

Michael A. White, Forrest Hoffman, William W. Hargrove, and Ramakrishna R. Nemani. A global framework for monitoring phenological responses to climate change. Geophys. Res. Lett., 32(4):L04705, February 2005. doi:10.1029/2004GL021961.

Yong Xue, Forrest M. Hoffman, and Dingsheng Liu. GeoComputation 2009. In Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M.A. Sloot, editors, Proceedings of the 9th International Conference on Computational Science (ICCS 2009), volume 5545 of Lecture Notes in Computer Science (LNCS), pages 345–348, Heidelberg, May 2009. Springer-Verlag. ISBN 978-3-642-01972-2. doi:10.1007/978-3-642-01973-9 38.